OpenStudy (anonymous):

How do you know if a limit is going to be zero when x is approaching infinity?

OpenStudy (bahrom7893):

if the denominator of a fraction is increasing at a faster rate than the numerator.

OpenStudy (anonymous):

what do you mean by increasing at a faster rate?

OpenStudy (bahrom7893):

for example: x/x^2 approaches 0 as x approaches infinity, because x approaches infinity slower than x^2

OpenStudy (bahrom7893):

Example: 1/1 = 1 2/4 = 0.5 3/9 = 0.333.... 4/16 = 0.25 5/25 = 0.2 6/36 = 0.166..

OpenStudy (bahrom7893):

See, the bottom increases faster than the top.. The example above was for x/x^2

OpenStudy (bahrom7893):

So if you keep going: 1,000,000/1,000,000,000,000 = 0.000001

OpenStudy (anonymous):

ohhhh i see. so if it is (e^x)/(x^2), the limit would be zero?

OpenStudy (bahrom7893):

no, the limit would be infinity. e^x increases faster than x^2

OpenStudy (anonymous):

oh whoops

OpenStudy (bahrom7893):

e=2.71 x=1 e^2 = something like 5 or 6 2^2 would be 4

OpenStudy (bahrom7893):

So the one you gave would go off to infinity.

OpenStudy (bahrom7893):

anything to the x increases faster than x to some constant power.

OpenStudy (bahrom7893):

as x goes to infinity.

OpenStudy (anonymous):

so lnx/(e^x) would have a limit of zero right

OpenStudy (bahrom7893):

yes

OpenStudy (anonymous):

ohhhh okay. thanksss